# An LLM-based Recommender System Environment

### v1

@@ Abstract
Reinforcement learning (RL) has gained popularity in the realm of recommender systems due to its ability to optimize long-term rewards and guide users in discovering relevant content. However, the successful implementation of RL in recommender systems is challenging because of several factors, including the limited availability of online data for training on-policy methods. This scarcity requires expensive human interaction for online model training. Furthermore, the development of effective evaluation frameworks that accurately reflect the quality of models remains a fundamental challenge in recommender systems. To address these challenges, we propose a comprehensive framework for synthetic environments that simulate human behavior by harnessing the capabilities of large language models (LLMs). We complement our framework with in-depth ablation studies and demonstrate its effectiveness with experiments on movie and book recommendations. By utilizing LLMs as synthetic users, this work introduces a modular and novel framework for training RL-based recommender systems. The software, including the RL environment, is publicly available. [^1] https://github.com/giorgiopiatti/SUBER
@@ 1 Introduction
In an age defined by the ubiquitous presence of digital platforms in both leisure and commerce, recommender systems have emerged as instrumental tools in guiding user choices. From Netflix tailoring movie suggestions to match cinematic tastes of users to Amazon presenting personalized lists of products to shoppers, recommendation systems are the engines driving enhanced user experiences and platform engagement (Steck et al., 2021; Agrawal et al., 2023). Reinforcement Learning (RL), with its principles rooted in learning by interaction, provides a compelling approach to dynamically and adaptively tailor recommendations. Recommender systems should take into account both short and long term rewards and direct the interests of users towards appropriate recommendations. An increasing body of research has investigated the use of RL in recommender systems (Ie et al., 2019b; Chen et al., 2019a; Liu et al., 2022; Afsar et al., 2022; Lin et al., 2023). Although promising, the use of RL for recommendation systems comes with its own set of challenges:
@@
Data Availability: RL algorithms require a significant amount of data from interactions with the environment to learn effective policies. However, in the case of recommender systems, users may quickly abandon the service if they receive random or irrelevant recommendations. This makes it impractical to collect the large amount of data needed to train an RL model without compromising the user experience (Zhang et al., 2016).
@@
Unknown user model: In RL, a reward function is crucial to allow the model to learn effectively. In the context of recommender systems, designing an appropriate synthetic reward function that accurately reflects user satisfaction or preferences can be challenging due to the complexity of modeling human behavior (Chen et al., 2019b; Shi et al., 2019).
@@
Model evaluation: A key challenge in recommender systems is the evaluation of models without directly interacting with real users, thus avoiding any potential negative impact on the user experience. On the other hand, evaluating on offline data does not guarantee good recommendation performance in the real world (Shani and Gunawardana, 2011; Garcin et al., 2014).
@@
In this work, we propose a “Simulated User Behavior for Recommender Systems” (SUBER), a novel framework for recommender systems to address the aforementioned challenges. SUBER is a framework for synthetic environments that use Large Language Models (LLM) at their core. SUBER leverages recent advances in LLMs to simulate human behavior (Park et al., 2023; Argyle et al., 2023). Furthermore, by training on large amounts of data, LLMs have obtained inherent knowledge about movies, books, and various other objects. These strengths, the ability to mimick human behavior coupled with vast knowledge about humanity, uniquely position LLMs as a powerful tool to simulate users in synthetic environments for recommender systems. Therefore, SUBER serves as a versatile playground, allowing researchers to experiment with different LLM configurations, fine-tune user specifications, and improve their RL strategies. Our contributions can be summarized as follows:
@@
  • We propose SUBER, a general framework to train and evaluate RL-based recommender systems. Our framework consists of a gym environment containing an LLM that is capable of simulating human behavior and rate recommended items accordingly.
@@
  • We conduct extensive ablation studies to examine the influence of the different components in our proposed framework. Moreover, we present findings across several families of LLMs and their effect on the performance of the environment, highlighting the effectiveness of various LLMs in replicating human choices for item recommendations.
@@
  • We experimentally assess our environment by evaluating both a movie recommendation setting and a book recommendation setting. We also make all code available as open-source.
@@ 2 Related work
Table 1: Comparison of simulation environments for RecSys. We list whether the user and item dataset are real or synthetic. Simulation Engine indicates the different approaches used. For the evaluation strategy, we distinguish between offline evaluation on the original dataset used to train the simulator, online testing on a platform, sanity checks, and case studies.
@@
Adversarial (Chen et al., 2019b) Real Real GAN Offline
VirtualTaobao (Shi et al., 2019) Real Real GAN Online
RL4RS (Wang et al., 2023a) Real Real Transformer Online
KuaiSim (Zhao et al., 2023) Real Real Transformer Offline
RecoGym (Rohde et al., 2018) Synthetic Synthetic Statistical modelling Sanity checks
RecoSim (Ie et al., 2019a) Synthetic Synthetic Statistical modelling Case studies
@@
Recommender systems. Platforms such as YouTube (Ie et al., 2019b; Chen et al., 2019a) and BytePlus (Liu et al., 2022) are two of many recent successful examples of training and evaluating recommender systems with online data. Traditional neural recommender systems and have been extensively researched in the past three decades (Goldberg et al., 1992; Su and Khoshgoftaar, 2009; Bobadilla et al., 2013; Shi et al., 2014; Lu et al., 2015; Zhang et al., 2019). In the area of RL for recommender systems, most of the work has been based mainly on training and evaluation based on offline datasets (Afsar et al., 2022; Lin et al., 2023). However, it presents significant challenges with respect to complexity and expense (Afsar et al., 2022). In contrast, offline evaluation takes place in a static and biased environment. To this end, many simulation environments for recommender systems have been developed (Rohde et al., 2018; Ie et al., 2019a; Chen et al., 2019b; Shi et al., 2019; Wang et al., 2023a; Zhao et al., 2023), and we present in Table 1 a qualitative comparison. Unlike previous approaches, our work leverages natural language by using LLMs to simulate user behavior. In addition, our framework is not dataset dependent, and therefore, the set of users and items are not restricted to specific domains.
Large Language Models. There have been significant recent advances in the field of LLMs (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023a, b; Jiang et al., 2023). LLMs can generate textual content that rivals the quality of human-generated text (Brown et al., 2020). Park et al. (2023) demonstrated how LLMs can be used to simulate human behavior. Furthermore, Argyle et al. (2023) suggests the use of language models as human surrogates within social science research. Contemporary work has also integrated LLMs into recommender systems. Kang et al. (2023) demonstrated that fine-tuned LLMs outperform traditional supervised methods in predicting user ratings with fewer training data, while Wang et al. (2023b) employed LLMs as a recommendation agent, showcasing their potential to improve recommender systems. Our work diverges from these approaches by using LLMs as simulation environments for item recommendation, in contrast to prior efforts that focused on training LLMs to function as the recommender system itself.
@@ 3 Framework
To address the aforementioned challenges of data availability, unknown user model, and model evaluation, we propose SUBER, an environment designed to simulate human behavior through the integration of LLMs. SUBER serves a dual purpose by generating synthetic data and harnessing the capabilities of LLMs to replicate the behavior of individuals with unknown patterns. Additionally, this dynamic environment can function as a model evaluation tool for recommender systems.
SUBER consists of an LLM component and three separate modules that contain multiple individual components. An overview of the overall structure is presented in Figure 1. The internal memory module of the environment contains two separate datasets, one for users and one for items. The environment also includes a preprocessing module that retrieves raw data from the memory module and transforms it to ensure compatibility with the LLM. Finally, a postprocessing component transforms the output produced by the LLM before returning it to the RL model.
The interaction with an RL model involves the following information flow: initially, the environment selects a user from memory, along with their interaction history (i.e., items and associated ratings) as the observation for the RL model. The RL model then recommends an item to the user as its action, with an action space equal to the number of items in the environment. The action and observation are subsequently processed through the preprocessing module, the LLM component, and the postprocessing module. Finally, the environment returns a reward corresponding to the postprocessed rating predicted by the LLM. We describe each module in more detail in the following sections.
Our environment is designed with easy accessibility and extensibility in mind. Therefore, we chose a modular approach and based the environment interface on the standardized API of Gymnasium (Towers et al., 2023). Different components can be modified at will, providing additional flexibility in future design choices.
@@ 3.1 Memory
We introduce the following notation. We define $U$ as the set of users and $I$ as the set of items. For every user-item pair $(u,i)\in U\times I$, we have a set $R_{u,i}$ that records all interactions between user $u$ and item $i$. Similarly, for every user $u$ we define with $R_{u}$ the set of all interactions with all movies, defined as follows:
@@
The memory module consists of three components: an item dataset, a user dataset, and a record of all interactions between users and items. This interaction history stores the set of interactions $R_{u,i}$ for each user-item pair $(u,i)$. Every interaction between the RL model and the environment produces a new interaction record between a user and an item, which is appended to the interaction history.
@@
3.2 Preprocessing
Item retrieval. As the RL model interacts with the environment, the interaction history grows. It may be challenging to extract relevant information from long histories, and increasing history length will likely exceed the context length of current LLMs (Park et al., 2023). To address this issue, we propose an item-retrieval component responsible for retrieving the most appropriate items for the current query from the interaction history of a user. Additionally, as user interests and preferences may evolve over time, relying solely on user features may not accurately capture current interests. Therefore, historical rating data are used to provide a more detailed depiction of their evolving preferences.
Prompting. The prompting component aggregates the information retrieved by the item retrieval component, creating a prompt that contains the necessary details for the LLM, including the user and query item data. The objective of this prompt is to enable the LLM to accurately predict the rating of the current query item. An example of such a prompt as part of an interaction example can be seen in Figure 2.
@@ 3.3 Postprocessing
Reward perturbation. The reward perturbation component introduces noise into the ratings generated by the LLM. This component functions as a simulation of “concept drift" for users (Žliobaitė et al., 2016). Concept drift refers to the notion that users may change their interests over time and are unlikely to maintain static preferences.
Reward shaping. Similarly to the reward perturbation component, reward shaping modifies the reward. However, unlike the perturbed reward which is added to the memory, the reward modified by the reward shaping component is returned directly to the RL model and is not stored in memory. The reward shaping module aims to reflect changes in the reward that are not related to a change in the preference of a user, such as spontaneous decisions or fleeting interests.
@@ 4 Experiments
To evaluate SUBER, we followed the approach of Rohde et al. (2018) and Ie et al. (2019a). We perform case studies and sanity checks, which we present in Section 4.3 and Section 5 . To achieve this, we implemented a movie recommendation and a book recommendation environment in our framework. In the following sections, we discuss our implementation and design choices for these environments. For the movie setting, we use rewards from 1 to 10, similar to TMDB  [^2] https://www.themoviedb.org/, while for the book setting we use rewards from 1 to 5, as found in the Amazon Reviews Dataset (Ni et al., 2019). For both environments, we created a dataset of synthetic users using Vicuna (Zheng et al., 2023) with Guidance (Lundberg, 2023). To generate user descriptions, we condition the LLM with information such as the age, liked and disliked genres, hobbies, and profession of the user (cf. LABEL:listing:prompt). We generate the user age by sampling from the age distribution in the United States (Bureau, 2022), while hobbies and professions were sampled from a precompiled list (cf. Appendix F).
@@
Listing 1: User generation. Black text shows the template, red marks the sampled information from external distributions, blue indicates the content generated by LLM.
For the movie dataset we use MovieLens (Harper and Konstan, 2015). For the book dataset, we used a subset of the Amazon Book Dataset. For more details, see Appendix A and Appendix B.

### v2

# SUBER: An RL Environment with Simulated Human Behavior for Recommender Systems

@@ Abstract
Reinforcement learning (RL) has gained popularity in the realm of recommender systems due to its ability to optimize long-term rewards and guide users in discovering relevant content. However, the successful implementation of RL in recommender systems is challenging because of several factors, including the limited availability of online data for training on-policy methods. This scarcity requires expensive human interaction for online model training. Furthermore, the development of effective evaluation frameworks that accurately reflect the quality of models remains a fundamental challenge in recommender systems. To address these challenges, we propose a comprehensive framework for synthetic environments that simulate human behavior by harnessing the capabilities of large language models (LLMs). We complement our framework with in-depth ablation studies and demonstrate its effectiveness with experiments on movie and book recommendations. Using LLMs as synthetic users, this work introduces a modular and novel framework to train RL-based recommender systems. The software, including the RL environment, is publicly available on https://github.com/SUBER-Team/SUBER.
[…]
[backgroundcolor=boxgray,linewidth=0pt,innerleftmargin=5pt,innerrightmargin=5pt,innertopmargin=5pt,innerbottommargin=5pt]graybox
[…]
1200
@@ 1 Introduction
In an age defined by the ubiquitous presence of digital platforms in both leisure and commerce, recommender systems have emerged as instrumental tools in guiding user choices. From Netflix tailoring movie suggestions to match the cinematic tastes of users to Amazon presenting personalized products lists to shoppers, recommendation systems are the engines driving enhanced user experiences and the engagement of the platform [44, 3].
Reinforcement Learning (RL), with its principles rooted in learning by interaction, provides a compelling approach to dynamically and adaptively tailor recommendations. Recommender systems should take into account both short- and long-term rewards and direct the interests of users towards appropriate recommendations. An increasing body of research has investigated the use of RL in recommender systems [18, 9, 23, 1, 22]. Although promising, the use of RL for recommendation systems comes with its own set of challenges:
@@
Data Availability: RL algorithms require a significant amount of data from interactions with the environment to learn effective policies. However, in the case of recommender systems, users may quickly abandon the service if they receive random or irrelevant recommendations. This makes it impractical to collect the large amount of data needed to train an RL model without compromising the user experience [53].
@@
Unknown user model: In RL, a reward function is crucial to allow the model to learn effectively. In the context of recommender systems, designing an appropriate synthetic reward function that accurately reflects user satisfaction or preferences can be challenging due to the complexity of modeling human behavior [10, 42].
@@
Model evaluation: A key challenge in recommender systems is the evaluation of models without directly interacting with real users, thus avoiding any potential negative impact on the user experience. On the other hand, evaluating on offline data does not guarantee good recommendation performance in the real world [41, 13].
@@
In this work, we propose a "Simulated User Behavior for Recommender Systems" (SUBER), a novel framework for recommender systems to address the aforementioned challenges. SUBER is a framework for synthetic environments that use Large Language Models (LLM) at its core. SUBER leverages recent advances in LLMs to simulate human behavior [31, 4]. Furthermore, by training on large amounts of data, LLMs have obtained inherent knowledge about movies, books, and various other objects. These strengths, the ability to mimick human behavior coupled with vast knowledge about humanity, uniquely position LLMs as a powerful tool to simulate users in synthetic environments for recommender systems. Therefore, SUBER serves as a versatile playground for researchers, allowing them to experiment with different LLM configurations, fine-tune user specifications, and improve their RL strategies. Our contributions can be summarized as follows:
@@
  • We introduce SUBER, a versatile framework for training and evaluating RL-based recommender systems. Our framework includes a gym environment with an LLM designed to simulate human behavior and rate recommended items accordingly.
@@
  • We conduct extensive ablation studies to assess the impact of each component in our framework. Moreover, we present findings across multiple LLM families, revealing their influence on the environment’s performance and highlighting their effectiveness in replicating human behavior for item recommendations.
@@
  • We experimentally validate our environment using both movie and book recommendation settings. Additionally, we have made all code available as open-source.
@@ 2 Related Work
Table 1: Comparison of simulation environments for recommender systems. We list whether the user and item datasets are real or synthetic. Simulation Engine indicates the different approaches used. For the evaluation strategy, we distinguish between offline evaluation in the original dataset used to train the simulator, online testing on a platform, sanity checks, and case studies.
@@
Adversarial [10] Real Real GAN Offline
VirtualTaobao [42] Real Real GAN Online
RL4RS [50] Real Real Transformer Online
KuaiSim [54] Real Real Transformer Offline
RecoGym [37] Synthetic Synthetic Statistical modelling Sanity checks
RecoSim [17] Synthetic Synthetic Statistical modelling Case studies
@@
RL for Recommender Systems.
Platforms such as YouTube [18, 9] and BytePlus [23] are two of many recent successful examples of training and evaluating recommender systems with online data. Traditional and neural recommender systems and have been extensively researched in the past three decades [15, 45, 5, 43, 24, 52]. However, since our work focuses on RL in recommender systems (RL4Rec), we limit the related work to this area of research. Although RL4Rec has been the subject of several studies, most of the work has been based primarily on training and evaluation based on offline datasets [1, 22]. As indicated by Afsar et al. [1], online assessment is the preferred approach for evaluation. However, it presents significant challenges with respect to complexity and expense. In contrast, offline evaluation takes place in a static and biased environment. Therefore, Afsar et al. call for the creation of a versatile simulator for RL4Rec similar in nature to OpenAI’s Gym for conventional RL tasks [6]. Additional challenges exist in the wider domain of RL, specifically regarding issues related to off-policy learning and offline policy evaluation, which become even more complex when incorporated into recommender systems [33, 14, 21].
Notable efforts have been made to address the limitations of offline learning in recommender systems. To this end, many simulation environments for recommender systems have been developed. Rohde et al. [37] presented RecoGym, a synthetic environment that addresses exploding variance by simulating user responses to different recommendation strategies. RecSim [17] is a customizable synthetic simulation platform that incorporates various assumptions about user preferences, item familiarity, user latent states and dynamics, and choice models. Chen et al. [10] proposed a generator that captures the underlying distribution of historical user interactions and learns to generate realistic interactions. Extending this idea, Shi et al. [42] proposed VirtualTaobao, a virtual shopping environment, and demonstrated the superiority of policies developed in this framework over traditional supervised techniques in real-world settings. Wang et al. [50] introduced the RL4RS dataset to address the lack of validated simulation environments and advanced evaluation methods in RL-based recommender system research. The dataset is collected from a NetEase game and anonymized through a three-step process. Zhao et al. [54] propose KuaiSim, a versatile environment that provides user feedback with multi-behavior and cross-session responses, supporting three tasks: request-level list-wise recommendation task, whole-session-level sequential recommendation task, and cross-session-level retention optimization task. Unlike previous approaches, our work leverages natural language by using LLMs to simulate user behavior. In addition, our framework is not dataset dependent, and therefore, the set of users and items are not restricted to specific domains.
Large Language Models.
There have been significant recent advances in the field of LLMs. These models are primarily based on the transformer architectures introduced by Vaswani et al. [49] and have continued to grow in size, capability, and performance. The Generative Pre-trained Transformer (GPT) series by OpenAI [7, 30] is one of the most notable developments in this area, demonstrating the immense potential and scalability of transformer-based models. The recent release of foundation language models such as Llama-1 and Llama-2 [46, 47], has democratized the access to these large LLMs. This has paved the way for the creation of instruction-following models such as Vicuna [55] and Mistral [19]. Meanwhile, numerous efforts have focused on optimizing the memory consumption and inference speed of LLMs. For example, GPTQ Frantar et al. [12] compressed the model parameters to 4 bits, allowing larger models to run on hardware with less memory and without significant loss of performance.
LLMs can generate textual content that rivals the quality of human-generated text [7]. However, their applications go beyond text generation. Park et al. [31] demonstrated how LLMs can be used to simulate human behavior. These simulated agents wake up, cook, go to work, make decisions, and reflect on past experiences in a believable manner. Furthermore, Argyle et al. [4] suggests using language models as surrogates for certain demographic groups within social science research. Their study demonstrates how conditioning GPT-3 on the socio-demographic backgrounds of real human subjects can accurately replicate response distributions among diverse human subgroups.
Contemporary work has also integrated LLMs into recommender systems. Kang et al. [20] demonstrated that fine-tuned LLMs outperform traditional supervised methods in predicting user ratings with less training data, while Wang et al. [51] employed LLMs as a recommendation agent, showcasing their potential to improve recommender systems. Both works show how LLMs can act as a good predictor of the ratings that a user would assign to an item. The authors further investigated whether LLMs can also be used as a recommender directly; they restricted their experiment to choosing an item from a list of 100 items. However, this task is still challenging for LLMs, as they must have knowledge of the entire set of possible items to be recommended. The limited context length does not allow one to provide a list of all possible items in the prompt to an LLM. Therefore, to date, the application of large language models (LLMs) as recommender systems has yet to exceed the performance of traditional recommender systems, which encompass both classical supervised algorithms and those based on reinforcement learning techniques. Our work diverges from these approaches by leveraging LLMs as simulation environments for item recommendation, in contrast to prior efforts that focused on training LLMs to function as the recommender system itself.
@@ 3 Framework
To address the aforementioned challenges of data availability, unknown user model, and model evaluation, we propose SUBER, an environment designed to simulate human behavior through the integration of LLMs. SUBER serves a dual purpose by generating synthetic data and harnessing the capabilities of LLMs to replicate the behavior of individuals with unknown patterns. Additionally, this dynamic environment can serve as a model evaluation tool for recommender systems.
SUBER consists of an LLM component and three separate modules that contain multiple individual components. An overview of the overall structure is presented in Figure 1. The internal memory module of the environment contains two separate datasets, one for users and one for items. The environment also includes a preprocessing module that retrieves raw data from the memory module and transforms it to ensure compatibility with the LLM. Finally, a post-processing component transforms the output produced by the LLM before returning it to the RL model.
The interaction with an RL model involves the following information flow: initially, the environment selects a user from memory, along with their interaction history (i.e., items and associated ratings) as the observation for the RL model. The RL model then recommends an item to the user as its action, with an action space equal to the number of items in the environment. The action and observation are subsequently processed through the preprocessing module, the LLM component, and the postprocessing module. Finally, the environment returns a reward corresponding to the post-processed rating predicted by the LLM. We describe each module in more detail in the following sections.
Our environment is designed with easy accessibility and extensibility in mind. Therefore, we chose a modular approach and based the environment interface on the Gymnasium standardized API [48]. Different components can be modified at will, providing additional flexibility in future design choices.
@@ 3.1 Memory
We introduce the following notation. We define $U$ as the set of users and $I$ as the set of items. For every pair of user-items $(u,i)\in U\times I$, we have a set $R_{u,i}$ that records all interactions between user $u$ and item $i$. Similarly, for every user $u$ we define with $R_{u}$ the set of all interactions with all items, defined as follows:
@@
The memory module consists of three components: an item dataset, a user dataset, and a record of all interactions between users and items. This interaction history stores the set of interactions $R_{u,i}$ for each pair of user-items $(u,i)$. Every interaction between the RL model and the environment produces a new interaction record between a user and an item, which is added to the interaction history.
3.2 Pre-processing
Item Retrieval.
As the RL model interacts with the environment, the history of the interaction increases. It may be challenging to extract relevant information from long histories, and the increasing duration of the history will probably exceed the context length of current LLMs [31]. To address this issue, we propose an item-retrieval component responsible for retrieving the most appropriate items for the current query from the interaction history of a user. Additionally, as user interests and preferences may evolve over time, relying solely on user features may not accurately capture current interests. Therefore, historical rating data are used to provide a more detailed depiction of their evolving preferences.
Prompting.
The prompting component aggregates the information retrieved by the item retrieval component, creating a prompt that contains the necessary details for the LLM, including the user and query item data. The objective of this prompt is to enable the LLM to accurately predict the rating of the current query item. An example of such a prompt as part of an interaction example can be seen in Figure 2.
@@ 3.3 Postprocessing
Reward Perturbation.
The reward perturbation component introduces noise into the ratings generated by the LLM. This component functions as a simulation of “concept drift" for users [56]. Concept drift refers to the notion that users may change their interests over time and are unlikely to maintain static preferences.
Reward Shaping.
Similarly to the reward perturbation component, reward shaping modifies the reward. However, unlike the perturbed reward which is added to the memory, the reward modified by the reward shaping component is returned directly to the RL model and is not stored in memory. The reward shaping module aims to reflect changes in the reward that are not related to a change in the preference of a user, such as spontaneous decisions or fleeting interests.
@@ 4 Experiments
To evaluate SUBER, we followed the approach of Rohde et al. [37] and Ie et al. [17]. We perform sanity checks and case studies, which we present in Section 4.2 and Section 4.4. To achieve this, we implemented a movie recommendation and a book recommendation environment in our framework. In the following sections, we discuss our implementation and design choices for these environments, as well as our ablation study and experiments. For the movie setting, we use rewards from 1 to 10, similar to TMDB  [^3] https://www.themoviedb.org/, while for the book setting we use rewards from 1 to 5, as found in the Amazon Reviews Dataset [28].
For both environments, we created a dataset of synthetic users using Vicuna [55] with Guidance [25]. To generate user descriptions, we condition the LLM with information such as the age, liked and disliked genres, hobbies, and profession of the user (cf. LABEL:listing:prompt). We generate the user age by sampling from the age distribution in the United States [8],
We randomly select a hobby and a profession from predefined lists (cf.  Appendix G). These hobby lists are divided into two categories: one tailored to children (aged 4-17) and another for adults (aged 18-75). Users not of working age are assigned the profession “student," while those of retirement age are categorized as “retired." For the movie dataset we use MovieLens (ml-latest-small) [16] and collect the respective movie features from TMDB. For the book dataset, we used a subset of the Amazon Book Dataset. For more details, see Appendix A and Appendix B.
@@
Listing 1: User generation and characteristic assignment process example by Vicuna with guidance. Black text shows the template and the instruction, red text marks the sampled information from external distributions, blue text indicates the content generated by LLM.